### v1

# Updating GUT-Scale Pole Higgs Inflation After ACT

Abstract
Abstract: We consider models of chaotic inflation driven by the real parts of a conjugate pair of Higgs superfields involved in the spontaneous breaking of a grand unification symmetry at a scale assuming its value within MSSM. We combine a superpotential, which is uniquely determined by applying a continuous $R$ symmetry, with two fractional shift-symmetric Kähler potentials introducing two free parameters $(p,N)$. The inflationary observables provide an excellent match to the recent ACT data for $1.355\leq p\leq 6.7$ and $6\cdot 10^{-5}\leq N\leq 0.7$. The attainment of inflation allows for subplanckian inflaton values and possibly detectable primordial gravitational waves with $(p,N)$ values of order unity. A solution to the $\mu$ problem of MSSM and baryogenesis via non-thermal leptogenesis can be also accommodated extending the superpotential of the model with suitable terms.
PACs numbers: 98.80.Cq, 04.50.Kd, 12.60.Jv, 04.65.+e
I Introduction
The announcement of Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT) act ; actin fuelled a plethora of works act0 ; act2 ; act4 ; gup ; oxf ; indi ; kina ; actlee ; rhc ; rhb ; rha ; act5 ; nmact ; maity ; reh8 ; actattr ; act1 ; actj ; yin ; actpal ; act3 ; act6 ; actellis ; acttamv ; ketov ; r2a ; r2b ; r2drees ; r2mans ; r2li ; heavy ; fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5 which revise many well-motivated inflationary models – for a review see Ref. review . This is because the combination of the aforementioned data with the cosmic microwave background (CMB measurements by Planck plin and BICEP/Keck (BK) bcp , together with the Dark Energy Spectroscopic Instrument (DESI) Baryon Acoustic Oscillation (BAO) results desi suggests a value of the (scalar) spectral index $n_{\rm s}$ substantially larger than the one indicated from the Planck data plin . Namely, the so-called P-ACT-LB-BK18 data entails actin
$n_{\rm s}=0.9743\pm 0.0068\>\>\mbox{and}\>\>r\leq 0.038$ (1)
[…]
In one recent paper actpole we propose a variant of pole inflation terada ; pole ; pole1 ; sor ; tmhi ; prl which assures compatibility with the $(n_{\rm s},r)$ values in Eq. (1). This goal becomes possible by introducing a particular class of rational Kähler potentials $K$ which include one extra parameter $p$ compared to the $\alpha$- ($N$- in our notation) attractor models alinde ; eno7 ; tmodel . Most notably, this achievement is independent of the type of the monomial superpotential $W$ selected leading, thereby, to the establishment of a new set of $(N,p)$ attractors. Working in the context of Supergravity (SUGRA), in our present investigation we adapt our proposal in Ref. actpole to a Grand Unified Theory (GUT) including the inflaton in a GUT-scale Higgs sector. In a such case, both $W$ and $K$ are consistent with a gauge symmetry rendering our scheme much more well-defined. Following the terminology of Ref. actpole we here realize Tp-Model inflation identifying the inflaton with the real part of a conjugate pair of Higgs superfields – cf. Ref. sor ; tmhi . We therefore could name our model Tp-Model Higgs Inflation (HI) or, for short, ppHI – for other works on GUT-scale HI see e.g. Ref. hi1 ; hi2 ; hi3 ; hi4 ; hi5 ; nmBL ; jhep ; lrcs ; ighi .
Apart from the compatibility with Eq. (1) ppHI offers high enough $r$ values which can be hopefully tested in the near future det . Nonetheless, it also allows us to explore consequences beyond the inflationary phenomenology. Namely, embedding ppHI into a BLB-L extension of Minimal Supersymmetric Standard Model (MSSM) we can accommodate a resolution of the $\mu$ problem, following the mechanism in Ref. dvali ; tmhi , and acceptable baryogenesis via non-thermal leptogenesis (nTL) dreeslept ; zhang ; lept consistently with the gravitino ($\widetilde{G}$) constraint brand ; kohri . The success of our post-inflationary scenario bounds from above the ratio $\mu/m_{3/2}$ – where $m_{3/2}$ is the $\widetilde{G}$ mass – and so it gives us the opportunity to link the high- to low-scale supersymmetry (SUSY).
Below, in Sec. II, we describe the building blocks ($W$ and $K$’s) of our proposal, we outline the derivation of the inflationary potential in Sec. III and test our model against observations in Sec. IV. Finally, we expose a post-inflationary completion in Sec. V and summarize our conclusions in Sec. VI. Unless otherwise stated, we use units where the reduced Planck scale $m_{\rm P}=2.43\cdot 10^{18}~{\mbox{\rm GeV}}$ is equal to unity.
II Model Set-up
The starting point of our approach is the superpotential known from the models of F-term hybrid inflation fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5
$W=\lambda S\left(\bar{\Phi}\Phi-M^{2}/2\right)$ (2)
which is uniquely determined, at renormalizable level, by a gauge symmetry $\mathbb{G}$ and a convenient continuous $R$ symmetry. Here, $\lambda$ and $M$ are two constants which can both be taken positive by field redefinitions; $S$ is a left-handed superfield, singlet under $\mathbb{G}$; $\bar{\Phi}$ and $\Phi$ is a pair of left-handed superfields belonging to non-trivial conjugate representations of $\mathbb{G}$, and reducing its rank by their vacuum expectation values (v.e.vs). Just for definiteness we restrict ourselves to $\mathbb{G}_{B-L}=\mathbb{G}_{\rm SM}\times U(1)_{B-L}$ which consists the simplest GUT beyond the MSSM – here $\mathbb{G}_{\rm SM}$ is the gauge group of MSSM and $B$ and $L$ denote the baryon and lepton number. With the specific choice of $\mathbb{G}$ the $U(1)_{B-L}$ and $R$ charges of the various superfields are
$(B-L)(S,\bar{\Phi},\Phi)=(0,-2,2)~~~\mbox{and}~~~R(S,\bar{\Phi},\Phi)=(2,0,0).$ (3)
At the SUSY limit $W$ leads to a BLB-L phase transition since the vacuum of the model lies at the direction
$\langle{S}\rangle\simeq 0~~\mbox{and}~~\langle{\Phi}\rangle=\langle{\bar{\Phi}}\rangle\simeq M/\sqrt{2}~~\mbox{for}~~M\ll 1,$ (4)
i.e., $U(1)_{B-L}$ is spontaneously broken via $\langle{\Phi}\rangle$ and $\langle{\bar{\Phi}}\rangle$.
To realize ppHI using $W$ in Eq. (2) – with $S$ stabilized at the origin and the inflaton included in the $\bar{\Phi}-\Phi$ system – a careful choice of $K$ is imperative as in Ref. sor ; jhep ; nmBL . The proposed here $K$ includes two contributions without mixing between $\bar{\Phi}-\Phi$ and $S$, i.e.,
$\widetilde{K}_{\ell\rm st}=\widetilde{K}_{\ell}+K_{\rm st}~~\mbox{with}~~\ell=1,2$ (5)
from which $K_{\rm st}$ successfully stabilizes $S$ along the inflationary path – see below – without invoking higher order terms. We adopt the form su11
$K_{\rm st}=N_{\rm st}\ln\left(1+{|S|^{2}/N_{\rm st}}\right)~~\mbox{with}~~0<N_{\rm st}<6$ (6)
which parameterizes the compact manifold $SU(2)/U(1)$ with curvature $2/N_{\rm st}$. On the other hand, $\widetilde{K}_{\ell}$ is devoted to the inflationary part and includes two contributions: a real one $K_{\ell}$ which determine the Kähler  metric for the inflaton sector and an holomorphic (and anti-holomorphic) part $K_{\rm sh}$ which provide $\widetilde{K}_{\ell}$ with a shift symmetry assuring $\widetilde{K}_{\ell}=0$ during ppHI – cf. Ref. actpole . Therefore, $\widetilde{K}_{\ell}$ has the structure
$\widetilde{K}_{\ell}=K_{\ell}+K_{\rm sh}~~\mbox{with}~~\ell=1,2.$ (7)
The key ingredient of our proposal is the selection of a totally fractional $K_{\ell}$ for $\ell=1$, which reads
$K_{1}=\frac{N}{\left(1-|\Phi|^{2}-|\bar{\Phi}|^{2}\right)^{p}}$ (8a)
defined for $N>0$ and $|\Phi|^{2}+|\bar{\Phi}|^{2}<1$. Identical results can be achieved if we select for $\ell=2$
$K_{2}=\frac{N}{2\left(1-2|\Phi|^{2}\right)^{p}}+\frac{N}{2\left(1-2|\bar{\Phi}|^{2}\right)^{p}}\\$ (8b)
defined for $N>0$, $|\Phi|<1/\sqrt{2}$ and $|\bar{\Phi}|<1/\sqrt{2}$. Although $K_{2}$ seems to be more contrived than $K_{1}$ – due to the appearance of the same factor in the numerators and exponent in the denominators – both $K$’s share the same symmetries: These are invariant under $\mathbb{G}$ and the interchange $\Phi\leftrightarrow\bar{\Phi}$. As regards the parameters $(p,N)$ we expect that $N$ is constrained observationally – as for the $N$ attractors alinde ; tmhi ; sor – and confine ourselves conservatively in the range $0.1\leq p\leq 10$.
On the other hand, $K_{\rm sh}$ is common for both $\widetilde{K}_{\ell}$ and takes the form
$K_{\rm sh}=-\frac{N}{2\left(1-2\bar{\Phi}\Phi\right)^{p}}-\frac{N}{2\left(1-2\bar{\Phi}^{*}\Phi^{*}\right)^{p}},$ (9)
where star denotes complex conjugation. It is obvious that $K_{\rm sh}$ let unchanged the Kähler metric by construction. Actually, the function in the denominators of $K_{\rm sh}$ is the unique possible quadratic function consistent with the gauge symmetry – cf. Eq. (2). Both $K$’s in Eq. (7) parameterize hyperbolic Kähler manifolds but without constant curvatures unlike the case of T-Model HI sor .
III Inflationary Potential
The appropriateness of the proposed $W$ and $\widetilde{K}_{\ell\rm st}$ in Eqs. (2) and (5) can be verified if we identify the inflationary potential and the canonically normalized inflaton with those introduced in Ref. actpole within SUGRA.
To embark on it, we focus on the lagrangian density for the complex scalar fields $z^{\alpha}=S,\Phi,\bar{\Phi}$ – denoted by the same superfield symbol – which can be written as linde1
${\cal L}=\sqrt{-\mathfrak{g}}\left(-\frac{1}{2}\mathcal{R}+K_{{\alpha}{\bar{\beta}}}g^{\mu\nu}D_{\mu}z^{\alpha}D_{\nu}z^{*{\bar{\beta}}}-V\right),$ (10)
where $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-), $\mathcal{R}$ is the Ricci scalar. The Kähler metric $K_{{\alpha}{\bar{\beta}}}$ and its inverse $K^{{\bar{\beta}}{\alpha}}$ are defined as follows
$K_{{\alpha}{\bar{\beta}}}={K_{,z^{\alpha}z^{*{\bar{\beta}}}}}>0\>\>\>\mbox{with}\>\>\>K^{{\bar{\beta}}{\alpha}}K_{{\alpha}\bar{\gamma}}=\delta^{\bar{\beta}}_{\bar{\gamma}},$ (11)
where $z^{\alpha}$ are complex scalar fields – the symbol $z^{\alpha}$ as subscript denotes derivation with respect to (w.r.t) $z^{\alpha}$ – and $D_{\mu}$ is the gauge covariant derivative. Also $V$ is the SUGRA potential which can be found in terms of $W$ and $K$’s in Eqs. (2) and (5) via the formula
$V=e^{K}\left(K^{{\alpha}{\bar{\beta}}}{\rm F}_{\alpha}{\rm F}^{*}_{\bar{\beta}}-3{|W|^{2}}\right)+\frac{g^{2}}{2}\mbox{$\sum_{a}$}{\rm D}_{a}{\rm D}_{a},$ (12a)
where the F and D terms are defined as follows
${\rm F}_{\alpha}=W_{,z^{\alpha}}+K_{,z^{\alpha}}W~~\mbox{and}~~{\rm D}_{a}=z^{\alpha}\left(T_{a}\right)^{\beta}_{\alpha}K_{\beta}$ (12b)
with $K_{{\alpha}}={K_{,z^{\alpha}}}$ and the summation is applied over the generators $T_{a}$ of $\mathbb{G}$. If we parameterize $\Phi,\bar{\Phi}$ and $S$ as follows
$\Phi={\phi e^{i{\theta}}}\cos{\theta_{\Phi}},~~\bar{\Phi}={\phi e^{i{\bar{\theta}}}}\sin{\theta_{\Phi}}~~\mbox{and}~~S=({s+i\bar{s}})/{\sqrt{2}}\,$, (13)
with $0\leq{\theta_{\Phi}}\leq\pi/2$, we can easily deduce from Eq. (12a) that a D-flat direction occurs at
$\langle{\bar{s}}\rangle_{\rm I}=\langle{s}\rangle_{\rm I}=\langle{{\theta}}\rangle_{\rm I}=\langle{{\bar{\theta}}}\rangle_{\rm I}=0\>\>\>\mbox{and}\>\>\>\langle{{\theta_{\Phi}}}\rangle_{\rm I}={\pi/4},$ (14)
where the symbol $\langle{Q}\rangle_{\rm I}$ means the value of the quantity $Q$ during ppHI. Since $\langle{K}\rangle_{\rm I}=0$ and $\langle{K_{SS^{*}}}\rangle_{\rm I}=1$ for both $\widetilde{K}_{\ell\rm st}$ in Eq. (5) the only surviving term from $V$ in Eq. (12a) is
$V_{\rm HI}=\langle{V}\rangle_{\rm I}=\langle{e^{K}K^{SS^{*}}\,|W_{,S}|^{2}}\rangle_{\rm I}\,={\lambda^{2}(\phi^{2}-M^{2})^{2}}/{4}$ (15)
which reduces to the quartic potential considered in Ref. actpole for $M\ll 1$ – see below.
To specify the canonically normalized inflaton, we note that, for both $\widetilde{K}_{\ell\rm st}$’s in Eq. (5), $K_{{\alpha}{\bar{\beta}}}$ along the configuration in Eq. (14) takes the form
$\langle{K_{{\alpha}{\bar{\beta}}}}\rangle_{\rm I}={\sf diag}\left(M_{\bar{\Phi}\Phi},\langle{K_{SS^{*}}}\rangle_{\rm I}\right)$ (16)
where the matrix $M_{\bar{\Phi}\Phi}$ is specified as follows
[…]
where $f_{\rm T}=1-\phi^{2}$ and $f_{p}=1+p\phi^{2}$. Upon diagonalization of $M_{\bar{\Phi}\Phi}$ for $K=\widetilde{K}_{\rm 1st}$ we find its eigenvalues which are
$\kappa_{+}=\kappa=pNf_{p}/f_{\rm T}^{p+2}~~\mbox{and}~~\kappa_{-}=pN/f_{\rm T}^{p+1},$ (18a)
from which $\kappa_{+}$ equals to the double eigenvalue $\kappa$ of $M_{\bar{\Phi}\Phi}$ for $K=\widetilde{K}_{\rm 2st}$. Inserting Eqs. (13) and (16) in the second term of the right-hand side of Eq. (10) we can define the canonically normalized fields which are denoted by hat. Combining Eqs. (17) and (18a) we can establish for both $\widetilde{K}_{\ell\rm st}$ in Eq. (5) a unique expression regarding the normalized inflaton $\widehat{\phi}$, i.e.,
${d\widehat{\phi}}/{d\phi}=J=\sqrt{2\kappa}~~\mbox{for}~~\phi<1.$ (19)
This result is identical with the $\phi-\widehat{\phi}$ relation obtained for $K=K_{1\rm T}$ (or $\widetilde{K}_{1\rm T}$) in Ref. actpole . For the remaining fields of the $\bar{\Phi}-\Phi$ system we find
$\hskip-8.5359pt\widehat{\theta}_{\pm}=\sqrt{\kappa_{\pm}}\phi\theta_{\pm},\>\>\widehat{\theta}_{\Phi}=\sqrt{2\kappa_{-}}\phi\left(\theta_{\Phi}-{\pi}/{4}\right)~\mbox{for $K=\widetilde{K}_{\rm 1st}$},$
$\hskip-8.5359pt\widehat{\theta}_{\pm}={\sqrt{\kappa}}\phi\theta_{\pm},\hskip 11.38092pt\widehat{\theta}_{\Phi}=\sqrt{2\kappa}\phi\left(\theta_{\Phi}-{\pi}/{4}\right)\hskip 11.38092pt\mbox{for $K=\widetilde{K}_{\rm 2st}$,}$
where ${\theta}_{\pm}=\left(\bar{\theta}\pm{\theta}\right)/\sqrt{2}$ – needless to say, $(s,\bar{s})$ are canonically normalized since $\langle{K_{SS^{*}}}\rangle_{\rm I}=1$. Note, in passing, that the spinors $\psi_{S}$ and $\psi_{\Phi\pm}$ associated with the superfields $S$ and $\Phi-\bar{\Phi}$ are normalized similarly, i.e., $\widehat{\psi}_{\Phi+}=\sqrt{\kappa}\psi_{\Phi+}$ with $\psi_{\Phi\pm}=(\psi_{\Phi}\pm\psi_{\bar{\Phi}})/\sqrt{2}$.
Table 1: Mass-squared spectrum for $K=\widetilde{K}_{\ell\rm st}$ ($\ell=1,2$) along the path in Eq. (14).
Fields Eigen Masses Squared
states $\widetilde{K}_{\rm 1st} K=K~$$\widetilde{K}_{\rm 2st}$
4 real $\widehat{\theta}_{+}$ $m_{\theta+}^{2}$ $6H_{\rm HI}^{2}$
scalars $\widehat{\theta}_{\Phi}$ $m_{\theta_{\Phi}}^{2}$ $M^{2}_{BL}+6H_{\rm HI}^{2}$
$s,\bar{s}$ $m_{S}^{2}$ $6\left(1/N_{\rm st}+2f_{\rm T}^{p+2}/pN\phi^{2}f_{p}\right)H_{\rm HI}^{2}$
1 gauge $A_{BL}$ $M_{BL}^{2}$ $2g^{2}pN\phi^{2}/f_{\rm T}^{p+1}$ $2g^{2}pN\phi^{2}f_{p}/f_{\rm T}^{p+2}$
boson
44 Weyl $\widehat{\psi}_{\pm}$ $m^{2}_{\psi\pm}$ $6f_{\rm T}^{p+2}H_{\rm HI}^{2}/pN\phi^{2}f_{p}$}
[…]
We can verify that the inflationary direction in Eq. (14) is stable w.r.t the fluctuations of the fields $\chi^{\alpha}=\theta_{+},\theta_{\Phi}$ and $S$. To this end, we construct the mass-squared spectrum of those scalars. Taking the limit $\phi\sim 1$ we find the expressions of the masses squared $m^{2}_{\chi^{\alpha}}$ arranged in Table 1, which approach rather well the exact expressions taken into account in our numerical computation. These expressions assist us to appreciate the role of $N_{\rm st}>0$ in retaining positive $m^{2}_{S}$. Moreover, $m^{2}_{\chi^{\alpha}}\gg H_{\rm HI}^{2}=V_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$ – where $\phi_{\star}$ and $\phi_{\rm f}$ are the values of $\phi$ when $k_{\star}=0.05/{\rm Mpc}$ crosses the horizon of ppHI and at the end of ppHI correspondingly. In Table 1 we display also the masses, $M_{BL}$, of the gauge boson $A_{BL}$ and the masses of the corresponding fermions with the eigenstate $\widehat{\psi}_{\pm}$ defined as $\widehat{\psi}_{\pm}=(\widehat{\psi}_{\Phi+}\pm{\psi}_{S})/\sqrt{2}$. Since $\mathbb{G}_{B-L}$ is broken during ppHI, no cosmic strings are formed after its termination.
The derived mass spectrum can be employed in order to find the one-loop radiative corrections, $\Delta V_{\rm HI}$, to $V_{\rm HI}$. Considering SUGRA as an effective theory with cutoff scale equal to $m_{\rm P}$, the well-known Coleman-Weinberg formula can be employed self-consistently taking into account only the masses which lie well below $m_{\rm P}$, i.e., all the masses arranged in Table 1 besides $M_{BL}$ and $m_{{\theta}_{\Phi}}$ – cf. Ref. nmBL ; jhep . The resulting $\Delta V_{\rm HI}$ lets intact our inflationary outputs, provided that the renormalization-group mass scale $\Lambda$, is determined by requiring $\Delta V_{\rm HI}(\phi_{\star})=0$ or $\Delta V_{\rm HI}(\phi_{\rm f})=0$. E.g., imposing the former condition for $N_{\rm st}=1$ we find $\Lambda\simeq(2.2-5.1)\cdot 10^{-5}$ for the allowed parameter space exposed below. Under these circumstances, our inflationary predictions can be exclusively reproduced by using $V_{\rm HI}$ in Eq. (15).
IV Inflationary Requirements
Applying the standard formulas quoted in Ref. actpole for $V_{\rm I}=V_{\rm HI}$, we can compute a number of observational quantities, which assist us to qualify our inflationary setting. Namely, we extract the number $N_{\star}$} of e-foldings that the scale $k_{\star}$ experiences during ppHI and the amplitude $A_{\rm s}$ of the power spectrum of the curvature perturbations generated by $\phi$ for $\phi=\phi_{\star}$. These observables must be compatible with the requirements actin
${N_{\star}}\simeq 61.5+\ln{V_{\rm HI}(\phi_{\star})^{\frac{1}{2}}\over V_{\rm HI}(\phi_{\rm f})^{\frac{1}{4}}}~~\mbox{and}~~A_{\rm s}^{\frac{1}{2}}\simeq 4.618\cdot 10^{-5},$ (20)
where we consider an equation-of-state parameter $w_{\rm rh}\simeq 1/3$ correspoding to quartic potential which is expected to approximate rather well $V_{\rm HI}$ for $\phi\ll 1$. The constraints above allows us to extract $\lambda$ and $\phi_{\star}$ for any given $p$, NN and $M$. However, $M$ is irrelevant for the inflationary observables, provided that it is confined to values much lower than $m_{\rm P}$. Motivated by the unification of the gauge coupling constants with the particle content of MSSM, we determine $M$ identifying the unification scale $M_{\rm GUT}\simeq 2/2.433\cdot 10^{-2}$ with the value of $M_{BL}$ – see Table 1 – at the SUSY vacuum in Eq. (4). Given that $\langle{f_{\rm T}}\rangle\simeq 1$ for $M\ll 1$, we obtain, for both $\widetilde{K}_{\ell\rm st}$ in Eq. (5)
$M\simeq{M_{\rm GUT}}/{g\sqrt{2Np}}$ (21)
with $g\simeq 0.7$ being the value of the GUT gauge coupling constant. The success of our inflationary setting is assured if
$M\lesssim m_{\rm P}~~\Rightarrow~N\gtrsim N_{\rm min}={M_{\rm GUT}^{2}}/{2g^{2}pm_{\rm P}^{2}}$ (22)
and so a lower bound on $N$ is applied. We are then left with only two free input parameters $(p,N)$ which can be constrained computing $n_{\rm s}$, its running, $a_{\rm s}$ and $r$. These quantities can be found approximately applying the analytic expressions displayed in Ref. actpole . Namely, we found
$n_{\rm s}\simeq 1-\frac{p+2}{(p+1){N_{\star}}},~~a_{\rm s}\simeq-\frac{p+2}{(p+1)N^{2}_{\star}}$ (23a)
which are indepedent of the exponent $n=4$ of $\phi$ in Eq. (15) whereas $r$ for $n=4$ is written as
$r\simeq{4(4^{p}pN)^{1/(p+1)}}/{(p+1)N_{\star}^{(p+2)/(p+1)}}.$ (23b)
The expressions above for $n_{\rm s}$ and $a_{\rm s}$ are independent from $N$ in the lowest order expansion in powers of $1/{N_{\star}}$. This feuture, however, is not totally kept as we show numerically in Fig. 1, where we compare our outputs against the P-ACT-LB-BK18 data actin in the $n_{\rm s}-r$ plane. We draw solid, dashed, dot-dashed and dotted lines for $p=1,2,5$ and $10$ respectively and show the variation of $N$ along each line. For the two latter $p$ values $N_{\rm min}$’s from Eq. (22) are also displayed. We clearly see that $n_{\rm s}$ increases with $p$. However, no upper bound on $p$ can be inferred from the upper bound on $n_{\rm s}$ in Eq. (1) and therefore we take $p=10$ as the maximal $p$ value. It is evident that the whole observationally favored range in the $n_{\rm s}-r$ plane is covered varying $p$ and $N$. As a bonus, $r$ values for $(p,N)$’s of order unity lie within the reach of planned experiments det aiming to discover primordial gravitational waves.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for various NN’s indicated on the curves. The conventions for the selected $p$ values are shown in the legends. The marginalized joint $68\%$ $95\%$] regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
Our plot in Fig. 1 reveals that $n_{\rm s}$ slightly decreases as $N$ increases and so we can obtain a lower bound on the pNp-N plane. The relevant allowed region is shown in Fig. 2. It is bounded by (i) the dashed line which originates the bound on $r$ in Eq. (1), (ii) the dot-dashed line which comes from the lower bound on $n_{\rm s}$ in Eq. (1) and (iii) the dotted line along which Eq. (22) is saturated. Fixing $n_{\rm s}$ to its central observational value in Eq. (1), we obtain the solid line along which we obtain the following allowed ranges
$1.355\lesssim p\lesssim 6.7~~~\mbox{and}~~~6\cdot 10^{-5}\lesssim N\lesssim 0.7$ (24)
with ${N_{\star}}\simeq(54.7-56.8)$, $a_{\rm s}\simeq-4.7\cdot 10^{-4}$ and $r\gtrsim 2.8\cdot 10^{-4}$. The obtained $|a_{\rm s}|$’s remain negligibly small being, thereby, consistent with its $95\%$ observationally allowed margin actin .
Note, finally, that our proposal is stabilized against corrections from higher order terms – e.g., $(\bar{\Phi}\Phi)^{l}$ with $l>1$ in Eq. (2) – since $\phi_{\star}<1$, as dictated by Eq. (19). The closer to unity $\phi_{\star}$ is chosen, the largest $N_{\star}$} is obtained. To quantify the relevant tuning we estimate $\Delta_{\star}=1-\phi_{\star}$ which ranges in the interval $(0.8-28)\%$ for the values in Eq. (24). The maximal $\Delta_{\star}$ values are obtained for the largest $p$’s and $N$’s indicating, thereby that naturalness prefers $N$ and $p$ values of order unity.
Fig. 2: Allowed (shaded) region as determined by Eq. (20) in the pNp-N plane. The conventions adopted for the various lines are also shown.
V Post-Inflationary Completion
Our inflationary setting can be easily embedded in a BLB-L extension of MSSM promoting to gauge the pre-existing global $U(1)_{B-L}$ – cf. Ref. nmBL ; tmhi . The terms of the total super- and Kähler potentials which control the coexistence of the inflationary and the MSSM sectors are
$\Delta W$ $=$ $\lambda_{\mu}S{H_{u}}{H_{d}}\ +\ h_{ijN}N^{c}_{i}L_{j}{H_{u}}+\lambda_{[N^{c}}i]\bar{\Phi}N^{c2}_{i};\hskip 19.91684pt$ (25a)
$\Delta K$ $=$ $\mbox{$\sum_{a=1}^{5}$}N_{\rm st}\ln\left(1+|X_{\alpha}|^{2}/N_{\rm st}\right)$ (25b)
with  $X_{\alpha}=N^{c}_{i},L_{j},{H_{u}}$ and $H_{d}$}. Here we adopt the notation and the BLB-L and $R$ charges of the various superfields as displayed in Table 1 of Ref. nmBL . For simplicity we use the same prefactors of $\ln$ in all terms of $\Delta K$.
We assume that $X_{\alpha}$ are stabilized at zero during ppHI and so the inflationary trajectory in Eq. (14) has to be supplemented by the condition
$\langle{{H_{u}}}\rangle_{\rm I}=\langle{{H_{d}}}\rangle_{\rm I}=\langle{\widetilde{N}^{c}_{i}}\rangle_{\rm I}=0,$ (26)
where $\widetilde{N}^{c}_{i}$ are the scalar superpartners of the right-handed neutrinos $N^{c}_{i}$. The stability of this path can be checked, parameterizing the complex fields above as we do for $S$ in Eq. (13) – note that we use small letters for the real and imaginary components. The relevant masses squared are listed in Table 2 where we see that $M^{2}_{iN^{c}}>0$ and $m^{2}_{h+}>0$ for any $\phi_{\rm f}<\phi<1$ – here $m_{h\pm}$ are the eigenvalues associated with the Higgs eigenstates
$h_{\pm}=(h_{u}\pm{h_{d}})/\sqrt{2}\>\>\>\mbox{and}\>\>\>{\bar{h}}_{\pm}=({\bar{h}}_{u}\pm{\bar{h}}_{d})/\sqrt{2}.$ (27)
On the other hand, the positivity of $m^{2}_{h-}$ dictates the establishment of the inequality – cf. Ref. tmhi :
$\lambda_{\mu}<\lambda(1+N_{\rm st})\phi_{\rm f}^{2}/2N_{\rm st}.$ (28)
We can verify numerically that the requirement above is fulfilled for $N_{\rm st}=1$ and $\lambda_{\mu}\lesssim 2\cdot 10^{-5}$. Despite this low upper bound, $\lambda_{\mu}$ can become compatible with $\mu/m_{3/2}\sim 1$ as we show in Sec. A, where the generation of the $\mu$ term of MSSM is analyzed. On the other hand, the implementation of nTL is explored in Sec. B and our final results, including the post-inflationary constraints, are given in Sec. C. Note in passing that due to the extra contributions of Table 2 $\Delta V_{\rm HI}$ increases and so $\Lambda$ should be readjusted. E.g., for $p=2$ we find $\Lambda\simeq(4-5)\cdot 10^{15}~{\mbox{\rm GeV}}$ – hereafter we restore units, i.e., we take $m_{\rm P}=2.433\cdot 10^{18}~{\mbox{\rm GeV}}$.
Table 2: Mass-squared spectrum of the non-inflaton sector along the path in Eqs. (14) and (26) for $K=\widetilde{K}_{\ell\rm st}+\Delta K$.
Fields Eigen- Masses Squared
states $\widetilde{K}_{\rm 1st}+\Delta K K=K~$$\widetilde{K}_{\rm 2st}+\Delta K$
10 Real $h_{\pm},{\bar{h}}_{\pm}$ $m_{h\pm}^{2}$ $3H_{\rm HI}^{2}\left(1+1/N_{\rm st}\pm{2\lambda_{\mu}}/{\lambda\phi^{2}}\right)$
Scalars $\tilde{\nu}^{c}_{i},\bar{\tilde{\nu}}^{c}_{i}$ $m_{i\tilde{\nu}^{c}}^{2}$ $3H_{\rm HI}^{2}\left(1+1/N_{\rm st}+8\lambda^{2}_{iN^{c}}\phi^{2}/\lambda^{2}\phi^{2}\right)$
33 Weyl $N_{i}^{c}$ $M_{{iN^{c}}}^{2}$ $24\lambda^{2}_{iN^{c}}H_{\rm HI}^{2}/\lambda^{2}\phi^{2}$
Spinors
A Generation of the $\mu$ Term of MSSM
The origin of the $\mu$ term of MSSM can be explained if we combine the terms of Eq. (2) with the first term in Eq. (25a) working at the level of SUSY. In fact, the total low energy potential is
$V_{\rm tot}=V_{\rm F}+V_{\rm soft},$ (29)
where the first term includes the SUSY F-term contributions into $V_{\rm tot}$ which read
$V_{\rm F}$ $=$ $K_{\rm SUSY}^{{\alpha}{\bar{\beta}}}(W+\Delta W)_{{\alpha}}(W+\Delta W)^{*}_{{\bar{\beta}}}$ (30)
$=$ $\lambda^{2}|\bar{\Phi}\Phi-M^{2}/2|^{2}+\lambda^{2}|S|^{2}\left(|\bar{\Phi}|^{2}+|\Phi|^{2}\right)/Np+\cdots$
where $K_{\rm SUSY}$ can be obtained by expanding $\widetilde{K}_{\ell\rm st}$ in Eq. (5) for low field values with result
$K_{\rm SUSY}=Np\left(|\Phi|^{2}+|\bar{\Phi}|^{2}-\bar{\Phi}\Phi-\bar{\Phi}^{*}\Phi^{*}\right)+|S|^{2}.$ (31)
Also ellipsis in Eq. (30) includes terms which vanish at the SUSY vacuum of Eq. (4) due to the zero v.e.vs of ${H_{u}},{H_{d}}$ and $\widetilde{N}^{c}_{i}$. On the other hand, the contribution into $V_{\rm tot}$ from the soft SUSY-breaking terms read
$V_{\rm soft}=\left(\lambda A_{\lambda}S\bar{\Phi}\Phi-\lambda{\rm a}_{S}M^{2}S/2+{\rm h.c.}\right)+m_{\gamma}^{2}\left|X^{\gamma}\right|^{2},$ (32)
where $X^{\gamma}=S,\Phi,\bar{\Phi},{H_{u}}$ and $H_{d}$}. Also $m_{\gamma},A_{\lambda}$ and ${\rm a}_{S}$ are soft SUSY-breaking mass parameters much lower than $M$. Rotating $S$ in the real axis by an appropriate $R$ transformation, choosing conveniently the phases of $A_{\lambda}$ and ${\rm a}_{S}$ – so as $V_{\rm tot}$ in Eq. (29) to be minimized – and substituting the $\Phi$ and $\bar{\Phi}$ values from Eq. (4), we get
$\langle{V_{\rm tot}(S)}\rangle=\lambda^{2}{M^{2}S^{2}}/{Np}-2\lambda{\rm a}_{\mu}m_{3/2}M^{2}S,$ (33a)
where $m_{3/2}$ is the $\widetilde{G}$ (gravitino) mass and
${\rm a}_{\mu}={(|A_{\lambda}|+|{\rm a}_{S}|)}/{2m_{3/2}}>0$ (33b)
is a parameter of order unity which parameterizes our ignorance about the dependence of $|A_{\lambda}|$ and $|{\rm a}_{S}|$ on $m_{3/2}$. The extermination condition of $\langle{V_{\rm tot}(S)}\rangle$ w.r.t $S$ shifts slightly $\langle{S}\rangle$ from zero in Eq. (4) – cf. Ref. dvali – as follows
${d}\langle{V_{\rm tot}(S)}\rangle/{dS}=0~~\Rightarrow~~\langle{S}\rangle=2Np\ {\rm a}_{\mu}m_{3/2}/{\lambda}.$ (33c)
Given the $\lambda-A_{\rm s}$ condition implied by the rightmost expression in Eq. (20), the derived $\langle{S}\rangle$ turns out to be several orders of magnitude below $m_{3/2}$ since $\lambda\sim 10^{-6}$. On the other hand, the generated $\mu$ parameter from the first term in Eq. (25a) is
$\mu=\lambda_{\mu}\langle{S}\rangle=2\lambda_{\mu}Np\ {\rm a}_{\mu}m_{3/2}/{\lambda}.$ (34)
It can be comparable to $m_{3/2}$ for $N\sim 1$, if $\lambda_{\mu}\sim 10^{-6}$. This prerequisite is welcome in view of Eq. (28) since it assures the stability of the path in Eq. (26) during ppHI.
B Non-Thermal Leptogenesis and $\widetilde{G}$ Constraint
Besides the generation of the $\mu$ term, $\Delta W$ in Eq. (25a) allows for the implementation of nTL via the direct decay of the inflaton into $N^{c}_{i}$. For the success of this process, however, one has to restrict the decay channel of the inflaton into $H_{u}$} and $H_{d}$} which opens via the first term of $\Delta W$. This implies constraints on $\lambda_{\mu}$ or, via Eq. (34), on the ratio $\mu/m_{3/2}$.
More explicitly, soon after the end of ppHI, the (canonically normalized) inflaton
$\widehat{\delta\phi}=\langle{J}\rangle\delta\phi\>\>\>\mbox{with}\>\>\>\delta\phi=\phi-M\>\>\>\mbox{and}\>\>\>\langle{J}\rangle\simeq\sqrt{2Np}$ (35)
acquires mass – reducing with increasing $N$ and $p$ – given by
$\widehat{m}_{\rm\delta\phi}=\left\langle V_{\rm HI,\widehat{\phi}\widehat{\phi}}\right\rangle^{1/2}\simeq{\lambda M}/{\sqrt{Np}}.$ (36)
Meanwhile, $\widehat{\delta\phi}$ settles into a phase of damped oscillations abound the minimum in Eq. (4) reheating the universe at a temperature reh ; reh1
$T_{\rm rh}=\left({90\over\pi^{2}g_{\rm rh*}}\right)^{1/4}\left({2(w_{\rm rh}-1)\over 5-3w_{\rm rh}}\widehat{\Gamma}_{\rm\delta\phi}m_{\rm P}\right)^{1/2}$ (37)
with $w_{\rm rh}\simeq 1/3$, $g_{\rm rh*}=228.75$ the MSSM effective number of relativistic degrees of freedom and $\widehat{\Gamma}_{\rm\delta\phi}$ the total decay width of $\widehat{\delta\phi}$
$\widehat{\Gamma}_{\rm\delta\phi}=\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}+\widehat{\Gamma}_{\delta\phi\rightarrow H},$ (38)
where the individual decay widths are
$\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ $=$ $({g_{iN^{c}}^{2}}\widehat{m}_{\rm\delta\phi}/{16\pi})\left(1-{4M_{[N^{c}}i]^{2}}/{\widehat{m}_{\rm\delta\phi}^{2}}\right)^{3/2};\hskip 14.22636pt$ (39a)
$\widehat{\Gamma}_{\delta\phi\rightarrow H}$ $=$ ${2}g_{H}^{2}\widehat{m}_{\rm\delta\phi}/{8\pi}$ (39b)
with the relevant coupling constants being
$g_{iN^{c}}=\sqrt{2}\lambda_{iN^{c}}/{\langle{J}\rangle}\>\>\>\>\mbox{and}\>\>\>\>g_{H}={\lambda_{\mu}}/{\sqrt{2}}.$ (39c)
Moreover, $M_{[N^{c}}i]$ are the Majorana masses of $N^{c}_{i}$ identified by the third term in Eq. (25a) as
$M_{[N^{c}}i]=\sqrt{2}\lambda_{iN^{c}}M~~\mbox{with}~~\lambda_{iN^{c}}\leq 4\pi$ (40)
to be consistent with the perturbative bound. The lagrangian terms which describe the decay channels above are similar to those given e.g. in Ref. nmBL .
For $M_{[N^{c}}i]\gg T_{\rm rh}$, the out-of-equilibrium decay of $N^{c}_{i}$ via the second term in Eq. (25a) generates a lepton-number-asymmetry yield which is partially converted through sphaleron effects into a yield of the observed baryon-asymmetry of the universe lept ; zhang
$Y_{B}=-0.35\cdot{5-3w_{\rm rh}\over 4(1-w_{\rm rh})}{T_{\rm rh}\over\widehat{m}_{\rm\delta\phi}}{\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}\over\widehat{\Gamma}_{\rm\delta\phi}}\varepsilon_{L}.$ (41)
The simplified formula above can be derived assuming normal hierarchy for $M_{[N^{c}}i]$ – i.e., $M_{[N^{c}}1]\ll M_{[N^{c}}2],M_{[N^{c}}3]$ – and that $\widehat{\delta\phi}$ decays via $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ in Eq. (39a) exclusively into $N^{c}_{1}$. Under these circumstances, we can obtain a maximal value for the lepton-number-asymmetry $\varepsilon_{L}$ which is sasa ; dreeslept
$\varepsilon_{L}=-\frac{3}{8\pi}\frac{m_{\nu_{\tau}}M_{[N^{c}}1]}{\langle{{H_{u}}}\rangle^{2}}\>\>\mbox{where}\>\>m_{\nu_{\tau}}=\sqrt{\Delta m^{2}_{\oplus}}\simeq 0.05\>{\rm eV}$ (42)
is the mass of heaviest light neutrino $\nu_{\tau}$ and $\Delta m^{2}_{\oplus}$ neutop the atmospheric neutrino mass-squared difference. Also, we set $\langle{{H_{u}}}\rangle=174~{\mbox{\rm GeV}}$ adopting the large $\tan\beta$ regime of MSSM. The validity of Eq. (41) requires the kinematic allowance of the decay $\widehat{\delta\phi}\rightarrow 2N^{c}_{1}$ and the avoidance of any erasure of the produced lepton asymmetry yield. These conditions can be fulfilled if we respectively demand
$\widehat{m}_{\rm\delta\phi}\geq 2M_{[N^{c}}1]\>\>\>\mbox{and}\>\>\>M_{[N^{c}}1]\gtrsim 10T_{\rm rh}.$ (43)
The result of Eq. (41) has to meet the observational one act
$Y_{B}=\left(8.75\pm 0.085\right)\cdot 10^{-11}~~\mbox{at $95\%$ c.l.}$ (44)
On the other hand, $T_{\rm rh}$ must be compatible with constraints on the $\widetilde{G}$ abundance, $Y_{3/2}$, at the onset of Big Bang nucleosynthesis (BBN), which is estimated to be kohri
$Y_{3/2}\simeq 1.9\cdot 10^{-13}~T_{\rm rh}/{\mbox{\rm EeV}},$ (45)
where we take into account only thermal production of $\widetilde{G}$ and assume that $\widetilde{G}$ is much heavier than the MSSM gauginos. On the other hand, $Y_{3/2}$ is bounded from above in order to avoid spoiling the success of the BBN. For the typical case where $\widetilde{G}$ decays with a tiny hadronic branching ratio, the bounds on $Y_{3/2}$ imply the following upper bounds on $T_{\rm rh}$ brand ; kohri
[…]
at $95\%$ c.l. The bounds above can be somehow relaxed in the case of a stable $\widetilde{G}$ – cf. Ref. fhi5 .
Fig. 3: Contours yielding the central $Y_{B}$ in Eq. (44) consistently with Eqs. (1), (20), (28), (43) and (46) in the $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and various $\mu/m_{3/2}$ values indicated on the curves – recall that $1~{\mbox{\rm EeV}}=10^{9}~{\mbox{\rm GeV}}$. The gray [light gray] segments of the lines are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$].
C Results
The compatibility of the restrictions in Eqs. (28), (43), (44) and (46) together with those from Eqs. (1) and (20) is illustrated in Fig. 3 where we present allowed curves in $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and $\mu/m_{3/2}=0.01,0.1,1$ and $3$. The chosen $p$ value is very close to the $p$ values which yields the central $n_{\rm s}$ in Eq. (1), as shown in Fig. 2, and the one for $N_{\rm st}$ assists us to overcome the constraint of Eq. (28) without extensive tuning. Moreover, the considered $\mu/m_{3/2}$ ratios are encountered in various recent studies klp ; gabit ; gmssm of the parameter space of MSSM with $m_{3/2}$ values comparable to those considered in Eq. (46). Despite the obtained $\mu$ value is quite low for $\mu/m_{3/2}=0.01$, it is in general still allowed by the LEP lep exclusion limit on the chargino mass.
The ratio $\mu/m_{3/2}$ controls the $\lambda_{\mu}$ value derived by Eq. (34) and via it the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow H}$ in Eq. (39b) into the $\widehat{\Gamma}_{\rm\delta\phi}$ – see Eq. (38). As expected, the required by Eq. (44) $M_{1N^{c}}$ values increase with $\mu/m_{3/2}$ in order to reinforce the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ – see Eq. (39a) – into $\widehat{\Gamma}_{\rm\delta\phi}$. The increase of $\mu/m_{3/2}$ is constrained directly by Eq. (28) and indirectly by Eq. (46) since it causes an increase of $T_{\rm rh}$ – see Eq. (37). The gray [light gray] segments of the lines displayed in Fig. 3 are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$]. As shown from the lower plot in Fig. 3 the lines corresponding to $\mu/m_{3/2}=0.01$ and $0.1$ terminate for low $N$ values due to the upper bound on $T_{\rm rh}$ in Eq. (46). On the other hand, for $\mu/m_{3/2}=1$ and $3$ this termination occurs due to the saturation of the bound on $\lambda_{\mu}$ in Eq. (28). The $\mu/m_{3/2}=3$ line is cut for $N=7.5$ due to the leftmost inequality in Eq. (43). The competition of these two aforementioned constraints causes a progressive reduction of the length of the allowed contours as $\mu/m_{3/2}$ increases beyond $3$ until they shrink to one point for $\mu/m_{3/2}\simeq 5.5$ corresponding to $N=4$, $M_{1N^{c}}=76~{\mbox{\rm EeV}}$ and $T_{\rm rh}=1.73~{\mbox{\rm EeV}}$. Along the curves of Fig. 3 we obtain
$0.035\lesssim\lambda_{\mu}/10^{-6}\lesssim 40~~\mbox{and}~~0.14\lesssim\widehat{m}_{\rm\delta\phi}/{\mbox{\rm ZeV}}\lesssim 32,$ (47)
where $1~{\mbox{\rm ZeV}}=10^{12}~{\mbox{\rm GeV}}$. The largest $\lambda_{\mu}$ is obtained for $\mu/m_{3/2}=1$ and $N=0.3$ whereas the lowest one for $\mu/m_{3/2}=0.01$ and $N=8$. On the other hand, $\widehat{m}_{\rm\delta\phi}$ decreases with increasing $N$ independently from $\mu/m_{3/2}$. Had we used larger $N_{\rm st}$ values most of the curves with $\mu/m_{3/2}>1$ would have been excluded. Therefore, our scheme primarily prefers the hierarchy $\mu\leq m_{3/2}$ which is defended by various analysis klp ; gabit ; gmssm of MSSM parameter space.
As a bottom line, nTL is a realistic possibility within our setting since it can be reconciled with the $\widetilde{G}$ constraint for $m_{3/2}\sim 10~{\mbox{\rm TeV}}$ and several MSSM versions.
VI Conclusions
Prompted by the P-ACT-LB-BK18 data we updated pole HI (i.e. Higgs inflation) sor – or T-Model HI tmhi – by extending our proposal in Ref. actpole . Namely, we adopted the super- and Kähler potentials $W$ and $K$ in Eqs. (2) and (5) which lead to the inflationary potential in Eq. (15) and the kinetic mixing in Eq. (19). Confining $(p,N)$ in the ranges shown in Fig. 2 we achieved a nice covering of the present data – see Fig. 1. Fixing $n_{\rm s}$ to its central value in Eq. (1) we obtained the clear predictions of Eq. (24) regarding our free parameters. As a consequence, our present framework drastically reduces the allowed $(p,N)$ domain found for a gauge-singlet inflaton in Ref. actpole . It is gratifying that the resulting $r$ values for $(p,N)$’s of order unity could be observationally accessible in the foreseeable future. We also specified a post-inflationary completion, based on the extra $W$ and $K$ terms in Eqs. (25a) and (25b), which offers a nice solution to the $\mu$ problem of MSSM and allows for baryogenesis via nTL with $M_{1N^{c}}$ in the range $(10-76){\mbox{\rm EeV}}$ provided that $\mu/m_{3/2}\leq 5.5$.
Since our main aim here was the introduction of ppHI, we opted to utilize its embedding in a GUT based on $\mathbb{G}_{B-L}$. Obviously our scheme can be realized, employing the same $W$ and $K$’s, within other SUSY GUTs too based on more structured gauge groups provided that $\Phi$ and $\bar{\Phi}$ consist a conjugate pair of Higgs superfields – cf. Ref. hi5 ; hi4 . Special treatment is required if the Higgs superfields belong to the adjoint representation of a GUT – cf. Ref. lrcs . Moreover, given that the term $\lambda M^{2}S/4$ of $W$ in Eq. (2) plays no role during ppHI, our scenario can be implemented by replacing it with $\lambda^{\prime}S^{3}$ and identifying $\Phi$ and $\bar{\Phi}$ with the electroweak Higgs doublets $H_{u}$ and $H_{d}$ of the next-to-MSSM linde1 . In the last case a tighter connection of the high- to the low-energy phenomenology is expected.
Acknowledgments
I would like to thank F. Koutroulis, A. Masiero and Q. Shafi for interesting discussions during the Corfu Summer Institute, where parts of this work were presented.